Prev Next

SW&IT Meeting Minutes 2020-07-06

Attendees: Carl, David, Igor, Matthieu, Stephen, Trey, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200629
LR20200629 - Bec, Matthieu 32.0
LR20200629 - Holmberg, Carl 32.0
LR20200629 - Edmeades, David 32.0
LR20200629 - Sola, Igor 0.0
LR20200629 - Hooper, Stephen 80.0
LR20200629 - Zhang, Yang 32.0
LR20200629 - Zhang, Xianyu 40.0
LR20200629 - Bryant, Trey 0.0

News from management

  • new evaluation forms, due this week

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • 802.1x—enabling network registration rolling out from July through August 10
  • Alfio trending #1 on hackernews!
  • LBC telemetry lagging

Team reports

MB

  • Last week
    • OC, Management
    • Recrutiment policy and working group
    • IT8132 MODS1 agw
    • TMS CoDR GMT report
    • Trey's induction
    • sciarch/tms meetings
  • This week
    • PE's
    • S/W review committee
    • CALAS
    • PB team reconnect
    • Trey's induction
    • Espresso SPIE review
    • DMS
      • websockets, AdSec, etc.
      • LBC stiched png previewfigure
      • max_bytes for ovms

SH

  • CentOS 8 finished testing, missing packages discussion on Slack
  • VPN
  • Statelite image
  • RPM for DMS SSL
  • Network issues on July 4th
  • DIMM computer issues
  • 507-1 lockiup
  • AGW5 a little, Xianyu a little
  • Wired 802.1X discussion with UITS
  • Wiki was open to public registration (Trey found it)
  • Performance eval
  • Meetings
  • Jenny's ROWG document

next:

  • VPN release for testing
  • prep for SSD work -- VPN, mountain firewall issues, etc.
  • CentOS 8 planning
  • ASC North mark

DE

  • VPN question from GB over the weekend; no clear detail or followup from him. Related to Carl's issue this morning?
  • JC finding out just how many dongles a new Mac needs. I thought he'd added an Ethernet-->USB-C but he didn't. He is taking care of it
  • LF computer arrived over the weekend, priority to get it going for her
  • Veeam quote arrived, $3k inital for 60 total licenses. I have asked for an estimate of our ongoing charges after the first year
  • Worked through the Jim Fog to fix his (self-inflicted?) Vault connection issue.
    • Discovered that he had been acting as point of contact for Hanna re: Vault stuff. I contacted her directly and told her to come to IT instead.
  • PIT logrotate/delay compression/move to Synology working as intended.
  • Worked with Trey to get him settled with accounts and onto his LBTO VM
  • Safety training
Specific tasks for this week:
  • LF computer
  • Mac Mini finish
  • DMS.mountain F31 re-kickstart with shiny new RPMs

YZ

  • Previous:
    • Some TMS work (wiki, software task, meeting).
    • Some AO-Cravat work (Strehl code revision, new OSCO datasets checkout).
    • Some MODS AGW troubleshooting.
  • Next:
    • TMS work (work on rough software version, review Andrew's emails/definitions, discuss with Heejoo a good time to experiment Multiline TCP server).
    • AO-Cravat (continue software work on Strehl code and utilities, plus new OSCO datasets analysis in more details).
    • DMS LBC telemetry lagging behind?
    • Performance evaluation (new form).
    • Safty training if time available.

CH

  • Last Week:
    • SupportScripts: replacing h5csv with dms api call.
    • IT7319, leap-second.list expiration still reported by AOSR, PMCR, and ECS. Yang and I restarted all three on Thursday via TCSGUI. Warnings from AOSR and PMCR stopped, ECS had not as of Thursday evening.
    • IT8132, MODS1 wfs cam init failure caused by faulty power supply which was in process of being permanently replaced late Thursday.
  • This Week:
    • AHS: LBTI alarm patch
    • SupportScripts: finish replacing h5csv with dms api call.
    • IT7319, check TCS logs for continued leap-second warnings.

TB

  • Previous:
    • Getting setup on a VM.
    • Making sure I can connect to the VPN.
    • Getting some familiarity with wiki.
  • Next:
    • Compliance training to complete.
    • Getting a start on the project?

XZ

  • Last week
    • Merge new4D to SOUL master
    • Test the merged new4D-SOUL on the simulator
    • Skip frame thresholds modification and test
    • Revert SX AdSec bump changes and test
    • AO operation support
    • Meetings
  • This week
    • Test the merged new4D-SOUL on telescope
    • Following and support SOUL operation
    • Power Backplane action items

IS

Last week:
  • PIT
    • PIT admin tool backend work. Have both ends, the react and the django talking to each other, basic stuff right now, only authentication and I'm starting to build an endpoint for the Call for Proposals (CfP) data. Then I'll have to work with Michelle to see what she wants to do with that and how she wants it. Also I'd like to prepare a presentation to the group on the technology, which I think we can all benefit from. The more the merrier.
  • Support
    • On Wednesday night I was called on an AGW issue. Nothing special happened, it was a matter of restarting processes but it took me a while. I didn't have the passwords ready, I got confused with the usernames used, etc. Quite stressful, fortunately it wasn't too late and the telescope was closed. So after I was done I realized that the information I needed to tackle this issue was pretty small. So I'm building a wikipage, but it could be also a google document where I just put the least information possible but that still works, so while in support of AGW, reading a few lines quickly give me 80% of what I need and the 20% that is not there I have links.
    • LUCI having the wrong mask issue in fits file after a cabinet exchange. This is not a new issue. I looked at that a while back but no resolution was found, except restarting the software. I'm talking to Dave T. and Barry and will work on this this week.

This week:
  • Safety Training
  • RPM
    • I'd like to start moving the ball on this.
  • LUCI
    • Look at mask issue #IT-??? (there is one, have to find it)
    • Partner with Barry/Dave on testing #IT7952?
    • work on configuration of LUCI WebIO, the HTTP GET api seems to be the same but on this device seems to be not consistent. Have to work a bit more on it. David E. provided support for network installation and backup configuration if needed.
  • PIT
    • If I have time I'd like to finish the CfP endpoint on the PIT admin tool work.
  • OT
    • Partner with Andrew to work on MODS XSLT
Topic revision: r13 - 06 Jul 2020, IgorSola
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback